Alfonso Inc. acquired 100 percent of the voting shares of BelAire Company on January 1, 2020. In exchange, Alfonso paid $263,500 in cash and issued 100,000 shares of its own $1 par value common stock. On this date, Alfonso’s stock had a fair value of $15 per share. The combination is a statutory merger with BelAire subsequently dissolved as a legal corporation. BelAire’s assets and liabilities are assigned to a new reporting unit.

The following shows fair values for the BelAire reporting unit for January 1, 2020 along with respective carrying amounts on December 31, 2021.

BelAire Reporting Unit Fair Values
1/1/20
Carrying Amounts
12/31/21
Cash $ 68,000 $ 41,000
Receivables 182,500 236,000
Inventory 219,000 251,000
Patents 371,500 467,000
Customer relationships 603,500 574,000
Equipment (net) 404,500 339,000
Goodwill ? 562,000
Accounts payable (123,500 ) (188,000 )
Long-term liabilities (524,000 ) (452,000 )

Note: Parentheses indicate a credit balance.

  1. Prepare Alfonso’s journal entry to record the assets acquired and the liabilities assumed in the BelAire merger on January 1, 2020. Note: Enter cash paid and cash received as two separate amounts.

  2. On December 31, 2021, Alfonso opts to forgo any goodwill impairment qualitative assessment and estimates that the total fair value of the entire BelAire reporting unit is $1,335,000. What amount of goodwill impairment, if any, should Alfonso recognize on its 2021 income statement?
Answer #1
Journal entries
Date General Journal Debit Credit
Jan 1 2020 Cash 68000
Receivables 182500
Inventory 219000
Patents 371500
Customer Relations 603500
Equipment (net) 404500
Goodwill 562000
Accounts payable 123500
Long term liabilities 524000
Cash 263500
common stock (Face value) 100000
Additional Paid in capital 1400000
Calculations
1 Good will
Particulars Amount Workings
Cash paid 263500
Fair value of shares issued 1500000 (100000*$15 per share)
Consideration transferred 1763500 (263500+1500000)
Fair value of net asset acquired 1201500 (68000+182500+219000+371500+603500+404500-123500-524000)
Good will 562000 (1763500-1201500)
2 Common stock 100000 (100000*$1 par value)
3 Additional Paid in capital 1400000 (68000+182500+219000+371500+603500+404500+562000-123500-524000-100000-263500)

...

Good will Impairment

Total Fair value of entire BelAire Unit= 1,335,000

Fair value reporting of net assets as oin 31/12/2021 = 41000+235000+251000+467000+574000+379000-188000-452000)

=1307000

Impaired fair value of goodwill = 1335000-1307000 =28000

Book Value of Goodwill = 562000

Good will impairment loss = (28000-562000) = $(534000)